OpenStudy (anonymous):

Can someone please help me. Create a word problem, it must be an interest problem, or a mixture problem. Your word problem must include the following: a word problem, the work to solve the word problem, and the solution .

OpenStudy (anonymous):

Heres my mixture problem. :) How many liters of 20% alcohol solution should be added to 40 liters of a 50% alcohol solution to make a 30% solution? Solution to problem. Let x be the quantity of the 20% alcohol solution to be added to the 40 liters of a 50% alcohol. Let y be the quantity of the final 30% solution. Hence x + 40 = y We shall now express mathematically that the quantity of alcohol in x liters plus the quantity of alcohol in the 40 liters is equal to the quantity of alcohol in y liters. But remember the alcohol is measured in percentage term. 20% x + 50% * 40 = 30% y Substitute y by x + 40 in the last equation to obtain. 20% x + 50% * 40 = 30% (x + 40) Change percentages into fractions. 20 x / 100 + 50 * 40 / 100= 30 x / 100 + 30 * 40 / 100 Mutliply all terms by 100 to simplify. 20 x + 50 * 40 = 30 x + 30 * 40 Solve for x. x = 80 liters 80 liters of 20% alcohol is be added to 40 liters of a 50% alcohol solution to make a 30% solution.

OpenStudy (anonymous):

I hope that makes sense :) Lol

OpenStudy (anonymous):

Actually it does make sense thank you so much.

OpenStudy (anonymous):

Haha okay good, lol no problem! :)